# The consistency strength of determinacy when all sets are universally Baire

### v2

@@ 4. The translation procedure
It is a well-known fact for countably complete extenders that they provide realizations, see for example [18, Lemma 10.64]. We will crucially use (a generalization of) this in the proof of Lemma 6.3 to obtain iterability for our translated structures.
@@
Now we can define our translation procedure. We aim to construct a premouse with strong cardinals $\kappa_{i}$, $i<\omega$, inside $\mathcal{W}$. The cardinals $\kappa_{i}$ are going to be the least ${<}\delta_{i+1}$-strong cardinals in $\mathcal{W}$. The idea is that extenders with critical point $\kappa\neq\kappa_{i}$ for all $i<\omega$ come from a standard fully backgrounded extender construction as in [8] (without the smallness assumption). In addition, we add extra extenders witnessing that $\kappa_{i}$ is strong for each $i<\omega$. These extenders with critical point $\kappa_{i}$ are generically countably complete. This will later allow us to prove that certain elementary substructures of the translated premouse are in fact sufficiently iterable (cf. Section 6.1) to show that the construction converges and the cardinals $\kappa_{i}$ are indeed fully strong101010“Fully strong” simply means “strong” but we decided to highlight this here as well as at some places below in order to avoid confusion with the statement that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong. (cf. Section 6.3).
@@
Let $\mathcal{W}$ be a translatable structure with Woodin cardinals $(\delta_{i}\mid i<\omega)$. Let $(\kappa_{i}\mid i<\omega)$ be a sequence of cardinals such that for each $i<\omega$, $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$ in $\mathcal{W}$. We say $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi\leq\Omega)$ is the mouse construction (or short $\mathcal{M}$-construction) in $\mathcal{W}$ for $\Omega\leq\operatorname{Ord}$ if it is given inductively as follows.
@@
    For $\xi+1<\Omega$, if $\mathcal{M}_{\xi}=(J_{\alpha}^{\vec{E}},\in,\vec{E},\emptyset)$ is a passive premouse, we define $\mathcal{N}_{\xi+1}$ as follows.
@@
      If there is an extender $F^{*}$ on the sequence of $\mathcal{W}$ and an extender $F$ over $\mathcal{M}_{\xi}$, both with critical point $\kappa\notin\{\kappa_{i}\mid i<\omega\}$ or $\kappa=\kappa_{i}$ for some $i<\omega$ with $\alpha<\delta_{i+1}$, such that $(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is a premouse and for some $\nu<\alpha$, $V_{\nu+\omega}^{\mathcal{W}}\subseteq\operatorname{Ult}(\mathcal{W},F^{*})$ and
@@
      If (2a) fails and there is a generically countably complete extender $F$ in $\mathcal{W}$ with critical point $\kappa_{i}$ for some $i<\omega$ such that $\delta_{i+1}\leq\alpha$ and $(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is a premouse, we let
@@
      If both (2a) and (2b) fail, we let
@@
    For $\xi+1<\Omega$, if $\mathcal{M}_{\xi}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is active, we let
@@
    If $\lambda\leq\Omega$ is a limit ordinal or $\lambda=\operatorname{Ord}$, we let $\mathcal{N}_{\lambda}=\mathcal{M}_{\lambda}$ be the unique passive premouse such that for all ordinals $\beta$, $\omega\beta\in\mathcal{N}_{\lambda}\cap\operatorname{Ord}$ iff $J_{\beta}^{\mathcal{N}_{\alpha}}$ is defined and eventually constant as $\alpha$ converges to $\lambda$, and for all such $\omega\beta\in\mathcal{N}_{\lambda}\cap\operatorname{Ord}$, $J_{\beta}^{\mathcal{N}_{\lambda}}$ is given by the eventual value of $J_{\beta}^{\mathcal{N}_{\alpha}}$ as $\alpha$ converges to $\lambda$.
@@
For the rest of this article, we let $\mathcal{M}$ be the result of an $\mathcal{M}$-construction up to $\Omega$ in the translatable structure $\mathcal{W}$ and argue that it converges, i.e., $\Omega=\operatorname{Ord}$, and that it has the desired properties.
@@ 5.1. Direct limit systems at $\kappa_{0}$
For simplicity, we start with a direct limit system at $\kappa_{0}$, the first cardinal that is strong up to $\delta_{1}$ in $\mathcal{W}$ [^12] Note that by a short argument in this setting $\kappa_{0}$ will also be the first cardinal that is strong up to $\delta_{1}$ in $\mathcal{M}$ since the fully backgrounded construction will add extenders witnessing this. The general case for $\kappa_{i}$, $i>0$, will be handled in the next section. We start working in $\mathcal{W}$. Let $\mathcal{M}^{0}_{\infty}$ be the direct limit of all $\Sigma_{0}$-iterates of $\mathcal{P}^{0}=Lp^{\infty}_{\omega}(\mathcal{W}|\delta_{0})$ that are countable in $\mathcal{W}[g]$ for some generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{0})$. By standard arguments this direct limit is well-defined. Let $\delta_{\infty}^{0}$ be the image of $\delta_{0}$ in $\mathcal{M}_{\infty}^{0}$ and write
$\mathcal{M}_{\infty}^{0,-}=\mathcal{M}_{\infty}^{0}|\delta_{\infty}^{0}.$
Note that $\mathcal{M}|\delta_{1}$ is equal to the result of a fully backgrounded extender construction in $\mathcal{W}|\delta_{1}$ (cf. [9, Lemma 2.21]). By the argument in [13, Lemma 2.14], $\kappa_{0}$ is a limit of Woodin cardinals in $\mathcal{M}$. We aim to identify $\mathcal{M}^{0}_{\infty}$ inside $\mathcal{M}$ as the limit of certain fully backgrounded constructions as follows. For any ordinal $\delta^{*}<\kappa_{0}$ that is a cutpoint in $\mathcal{M}|\kappa_{0}$, let
@@
be the result of a fully backgrounded $L[E]$-construction over $\emptyset$ in $\mathcal{M}|\delta_{1}$ using extenders with critical point above $\delta^{*}$, in other words, $\mathcal{L}_{\eta^{*}}=(L[E]^{\geq\delta^{*}}(\emptyset))^{\mathcal{M}|\delta_{1}}$. Let $\gamma\leq\eta^{*}$ be the least ordinal such that
@@
As $\kappa_{0}$ is a limit of Woodin cardinals in $\mathcal{M}$ (cf. [13, Lemma 2.14]) and the least ${<}\delta_{1}$-strong cardinal in $\mathcal{W}$, there are such Woodin cardinals in $(Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma}))^{\mathcal{W}}$ cofinally below $\kappa_{0}$. In particular, for any cutpoint $\delta^{*}<\kappa_{0}$ in $\mathcal{M}$ and $\gamma\leq\eta^{*}$ as above, $\mathcal{L}_{\gamma}\cap\operatorname{Ord}<\kappa_{0}$. We start with arguing that these models $\mathcal{N}_{\delta^{*}}$ and $\mathcal{N}_{\delta^{*}}^{+}$ are in fact inside $\mathcal{M}$. This is a consequence of the following more general lemma, see also [13, Lemma 2.12].
@@
The function $X\mapsto(Lp^{\infty}(X))^{\mathcal{W}}$ for $X\in\mathcal{M}|\kappa_{0}$ is definable over $\mathcal{M}|\delta_{1}$.
@@
Let $\delta^{*}<\kappa_{0}$ be a cutpoint in $\mathcal{M}|\delta_{1}$ with $X\cap\operatorname{Ord}<\delta^{*}$ and let $\mathcal{L}_{\eta^{*}}=(L[E]^{\geq\delta^{*}}(X))^{\mathcal{M}|\delta_{1}}$ be the result of a fully backgrounded $L[E]$-construction over $X$ in $\mathcal{M}|\delta_{1}$ using extenders with critical point above $\delta^{*}$. Let
@@
Let $\eta>X\cap\operatorname{Ord}$ be a cardinal in $\mathcal{M}|\kappa_{0}$. Then
@@
The direction $\mathcal{O}^{\mathcal{L}_{\eta^{*}}}_{\eta}\unlhd(Lp^{\infty}(\mathcal{L}_{\eta}))^{\mathcal{W}}$ is clear. For the other direction, fix an arbitrary sound $\mathcal{N}\unlhd Lp^{\infty}(\mathcal{L}_{\eta})$ in $\mathcal{W}$ with $\mathcal{L}_{\eta}\unlhd\mathcal{N}$ and $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$. Let $E$ be the index-least extender on the extender sequence of $\mathcal{L}_{\eta^{*}}$ such that $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ is a cutpoint in $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$. It suffices to show that $\mathcal{N}\unlhd\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$. Note that both $\mathcal{N}$ and $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ are fully iterable in $\mathcal{W}$ above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ as the extenders used to construct $\mathcal{L}_{\eta^{*}}$ are fully backgrounded by extenders in $\mathcal{M}|\delta_{1}$ which are themselves fully backgrounded by extenders in $\mathcal{W}$. This uses the observation that the extenders in $\mathcal{M}$ with index below $\delta_{1}$ are all added by the fully backgrounded part of the $\mathcal{M}$-construction.
By universality of fully backgrounded constructions (cf. [12, Lemma 2.13]), $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ wins the coiteration with $\mathcal{N}$. As $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$ and the coiteration is above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$, the $\mathcal{N}$-side cannot move as otherwise it would have to drop. Suppose the $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$-side moves to an iterate $\mathcal{W}^{*}$ of $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ such that $\mathcal{N}\unlhd\mathcal{W}^{*}$. If the iteration from $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ to $\mathcal{W}^{*}$ drops on the main branch, $\mathcal{W}^{*}$ is not fully sound, so $\mathcal{N}\lhd\mathcal{W}^{*}$. As $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$, $\mathcal{N}$ is in fact a subset of $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ in $\mathcal{W}^{*}$. But iterations above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ do not add any new subsets of $\mathcal{L}_{\eta}\cap\operatorname{Ord}$, thus $\mathcal{N}\unlhd\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$, as desired. If the iteration from $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ to $\mathcal{W}^{*}$ does not drop on the main branch, $\mathcal{W}^{*}\cap\operatorname{Ord}=\mathcal{L}_{\eta^{*}}\cap\operatorname{Ord}=\delta_{1}$ and again $\mathcal{N}\lhd\mathcal{W}^{*}$. So we can finish the argument as in the dropping case. ∎
The lemma now follows from the next claim, which is proved like Claim 2.13 in the proof of [13, Lemma 2.12]. It uses Clause (6) in Definition 3.20 and generic interpretability of the involved iteration strategies with branch condensation.
@@
For any $X\in\mathcal{M}|\kappa_{0}$, $\mathcal{N}\unlhd(Lp^{\infty}(X))^{\mathcal{W}}$ if, and only if, there is some $\mathcal{R}\unlhd\mathcal{O}^{\mathcal{L}_{\eta^{*}}}_{\eta}$ such that
@@
We start internalizing $\mathcal{M}_{\infty}^{0}$ into $\mathcal{M}$ with the following observation, that still takes place in $\mathcal{W}$.
@@
Let $\delta^{*}<\kappa_{0}$ be a cutpoint in $\mathcal{M}$. Then, in $\mathcal{W}$, $\mathcal{N}_{\delta^{*}}^{+}$ is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$ and the iteration does not drop on the main branch.
@@
Work in $\mathcal{W}$ and compare $\mathcal{M}|\delta_{0}$ against the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ giving rise to $\mathcal{N}_{\delta^{*}}$. By stationarity and universality of fully backgrounded constructions [^13] Cf. [12, Lemmas 2.11 and 2.13], note that these results also hold in our setting where $\mathcal{N}_{\delta^{*}}$ is the result of a fully backgrounded construction in $\mathcal{M}|\delta_{1}$, that itself is an initial segment of the result of a fully backgrounded construction in $\mathcal{W}$., $\mathcal{M}|\delta_{0}$ iterates via $\Sigma_{0}$ without drops on the main branch to some level of the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$. As $\mathcal{N}_{\delta^{*}}^{+}=Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma})$ for $\gamma$ the least ordinal such that $\mathcal{L}_{\gamma}\cap\operatorname{Ord}$ is Woodin in $Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma})$, it follows by super fullness preservation of $\Sigma_{0}$ that in fact $\mathcal{P}^{0}=Lp_{\omega}^{\infty}(\mathcal{W}|\delta_{0})$ iterates via $\Sigma_{0}$ to $\mathcal{N}_{\delta^{*}}^{+}$. ∎

### v3

@@ 4. The translation procedure
It is a well-known fact for countably complete extenders that they provide realizations, see for example [19, Lemma 10.64]. We will crucially use (a generalization of) this in the proof of Lemma 6.3 to obtain iterability for our translated structures.
@@
Now we can define our translation procedure. We aim to construct a premouse with strong cardinals $\kappa_{i}$, $i<\omega$, inside $\mathcal{W}$. The cardinals $\kappa_{i}$ are going to be the least ${<}\delta_{i+1}$-strong cardinals in $\mathcal{W}$. The idea is that extenders with critical point $\kappa\neq\kappa_{i}$ for all $i<\omega$ come from a standard fully backgrounded extender construction as in [8] (without the smallness assumption). In addition, we add extra extenders witnessing that $\kappa_{i}$ is strong for each $i<\omega$. These extenders with critical point $\kappa_{i}$ are still fully backgrounded in case they are added in the construction below the next Woodin cardinal $\delta_{i+1}$. These extenders will not overlap $\delta_{i+1}$. At levels of the construction past $\delta_{i+1}$ all extenders that are added with critical point $\kappa_{i}$ are generically countably complete (and these extenders will overlap $\delta_{i+1}$). This will later allow us to prove that certain elementary substructures of the translated premouse are in fact sufficiently iterable (cf. Section 6.1) to show that the construction converges and the cardinals $\kappa_{i}$ are indeed fully strong131313“Fully strong” simply means “strong” but we decided to highlight this here as well as at some places below in order to avoid confusion with the statement that $\kappa_{i}$ is ${<}\delta_{i+1}$-strong. (cf. Section 6.3).
@@
Let $\mathcal{W}$ be a translatable structure with Woodin cardinals $(\delta_{i}\mid i<\omega)$. Let $(\kappa_{i}\mid i<\omega)$ be a sequence of cardinals such that for each $i<\omega$, $\kappa_{i}$ is the least ${<}\delta_{i+1}$-strong cardinal above $\delta_{i}$ in $\mathcal{W}$. We say $(\mathcal{M}_{\xi},\mathcal{N}_{\xi}\mid\xi\leq\Xi)$ is the mouse construction (or short $\mathcal{M}$-construction) in $\mathcal{W}$ for $\Xi\leq\operatorname{Ord}$ if it is given inductively as follows.
@@
    For $\xi+1<\Xi$, if $\mathcal{M}_{\xi}=(J_{\alpha}^{\vec{E}},\in,\vec{E},\emptyset)$ is a passive premouse, we define $\mathcal{N}_{\xi+1}$ as follows.
@@
      If there is an extender $F^{*}$ on the sequence of $\mathcal{W}$ and an extender $F$ over $\mathcal{M}_{\xi}$, both with critical point $\kappa\notin\{\kappa_{i}\mid i<\omega\}$, [^14] In this case, we just follow the standard fully backgrounded construction in [8] without the smallness assumption. such that $(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is a premouse and for some $\nu<\alpha$, $V_{\nu+\omega}^{\mathcal{W}}\subseteq\operatorname{Ult}(\mathcal{W},F^{*})$ and
@@
      If (2a) fails and there is an extender $F^{*}$ on the sequence of $\mathcal{W}$ and an extender $F$ over $\mathcal{M}_{\xi}$, both with critical point $\kappa=\kappa_{i}$ for some $i<\omega$ with $\alpha<\delta_{i+1}$, [^16] In this case, we still follow the fully backgrounded construction as the new extender will not overlap any of the Woodin cardinals $\delta_{i}$. such that $(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is a premouse and for some $\nu<\alpha$, $V_{\nu+\omega}^{\mathcal{W}}\subseteq\operatorname{Ult}(\mathcal{W},F^{*})$ and
      we let
@@
      If (2a) and (2b) fail and there is a generically countably complete extender $F$ in $\mathcal{W}$ with critical point $\kappa_{i}$ for some $i<\omega$ such that $\alpha>\delta_{i+1}$ and $(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is a premouse, we let
      $\mathcal{N}_{\xi+1}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F).$
    4. (d)
      If (2a), (2b) and (2c) fail, we let
@@
    For $\xi+1<\Xi$, if $\mathcal{M}_{\xi}=(J_{\alpha}^{\vec{E}},\in,\vec{E},F)$ is active, we let
@@
    If $\lambda\leq\Xi$ is a limit ordinal or $\lambda=\operatorname{Ord}$, we let $\mathcal{N}_{\lambda}=\mathcal{M}_{\lambda}$ be the unique passive premouse such that for all ordinals $\beta$, $\omega\beta\in\mathcal{N}_{\lambda}\cap\operatorname{Ord}$ iff $J_{\beta}^{\mathcal{N}_{\alpha}}$ is defined and eventually constant as $\alpha$ converges to $\lambda$, and for all such $\omega\beta\in\mathcal{N}_{\lambda}\cap\operatorname{Ord}$, $J_{\beta}^{\mathcal{N}_{\lambda}}$ is given by the eventual value of $J_{\beta}^{\mathcal{N}_{\alpha}}$ as $\alpha$ converges to $\lambda$.
@@
Recall that we fixed a translatable structure $\mathcal{W}$. The purpose of the rest of this paper is to show that an $\mathcal{M}$-construction inside a translatable structure produces the desired output, i.e.,
  1. (a)
    it results in a proper class model $\mathcal{M}$ (i.e., $\Xi=\operatorname{Ord}$),
  2. (b)
    in $\mathcal{M}$, all $\kappa_{i}$ for $i<\omega$ are strong cardinals.
That all $\delta_{i}$ for $i<\omega$ are Woodin cardinals in $\mathcal{M}$ follows from the arguments in [8].
The argument in the rest of this paper is organized inductively on $i<\omega$. That means, at stage $i$ of the argument we inductively assume that if $\mathcal{M}^{(i)}$ denotes the result of an $\mathcal{M}$-construction where Clause (2c) is restricted to extenders with critical point in $\{\kappa_{j}\mid j<i\}$,
$(\mathsf{IH}.1)_{i}$:
for every $j<i$ and every $\bar{\mathcal{W}}$ the transitive collapse of a hull of $\mathcal{W}|\Omega$ for some sufficiently large $\Omega$ such that $\mathcal{W}|\kappa_{j}\subseteq\bar{\mathcal{W}}$, if
$\pi\colon\bar{\mathcal{W}}\rightarrow\mathcal{W}$
denotes the uncollapse embedding, $\bar{\mathcal{M}}$ the collapse of the result of an $\mathcal{M}$-construction where Clause (2c) is restricted to extenders with critical point in $\{\kappa_{j}\mid j<i\}$ in $\mathcal{W}|\Omega$ and $(\bar{\kappa}_{l},\bar{\delta}_{l}\mid l<\omega)$ the preimages of $(\kappa_{l},\delta_{l}\mid l<\omega)$, then
[…]
with respect to extenders with critical point above $\kappa_{j}$.
By running the proofs of Lemmas 6.5 and 6.6 this allows us to inductively assume that
$(\mathsf{IH}.2)_{i}$:
the construction of $\mathcal{M}^{(i)}$ does not break down, and
$(\mathsf{IH}.3)_{i}$:
all $\kappa_{j}$ for $j<i$ are strong in $\mathcal{M}^{(i)}$.
For each $i<\omega$, write $(\mathsf{IH})_{i}$ for the conjuction of $(\mathsf{IH}.1)_{i}$, $(\mathsf{IH}.2)_{i}$, and $(\mathsf{IH}.3)_{i}$.
@@ 5.1. Direct limit systems at $\kappa_{0}$
For simplicity, we start with a direct limit system at $\kappa_{0}$, the first cardinal that is strong up to $\delta_{1}$ in $\mathcal{W}$ [^17] Note that by a short argument due to Farmer Schlutzenberg in this setting $\kappa_{0}$ will also be the first cardinal that is strong up to $\delta_{1}$ in $\mathcal{M}^{(0)}$ since the fully backgrounded construction will add extenders witnessing this. The general case for $\kappa_{i}$, $i>0$, will be handled in the next section. For this case with $i=0$ we do not require any inductive hypothesis (note that $(\mathsf{IH})_{0}$ is an empty condition). Recall, that $\mathcal{M}^{(0)}$ denotes the result of an $\mathcal{M}$-construction where Clause (2c) is restricted to extenders with critical point $\kappa_{0}$. Note that $\mathcal{M}^{(0)}$ is not necessarily a proper class model at this point. Showing that the construction does not break down and $\mathcal{M}^{(0)}$ is indeed a proper class model is one of the applications of the results in this subsection.
We start working in $\mathcal{W}$. Let $\mathcal{P}^{0}_{\infty}$ be the direct limit of all $\Sigma_{0}$-iterates of $\mathcal{P}^{0}=Lp^{\infty}_{\omega}(\mathcal{W}|\delta_{0})$ that are countable in $\mathcal{W}[g]$ for some generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{0})$. By standard arguments this direct limit is well-defined and well-founded, see, for example, [31]. Let $\delta_{\infty}^{0}$ be the image of $\delta_{0}$ in $\mathcal{P}_{\infty}^{0}$ and write
$\mathcal{P}_{\infty}^{0,-}=\mathcal{P}_{\infty}^{0}|\delta_{\infty}^{0}.$
Note that $\mathcal{M}^{(0)}|\delta_{1}=\mathcal{M}|\delta_{1}$ is equal to the result of a fully backgrounded extender construction in $\mathcal{W}|\delta_{1}$ (cf. [9, Lemma 2.21]). Therefore, the $\mathcal{M}$-construction as well as the construction of $\mathcal{M}^{(0)}$ succeed up to $\delta_{1}$ and $\mathcal{M}^{(0)}|\delta_{1}=\mathcal{M}|\delta_{1}$ are well-defined. Also, if both constructions succeed far enough, $\mathcal{M}^{(0)}|\delta_{2}=\mathcal{M}|\delta_{2}$. By the argument in [14, Lemma 2.14], $\kappa_{0}$ is a limit of Woodin cardinals in $\mathcal{M}^{(0)}$. We aim to identify $\mathcal{P}^{0}_{\infty}$ inside $\mathcal{M}^{(0)}$ as a direct limit of models such that densely many of these models can be obtained via certain fully backgrounded constructions as follows. For any ordinal $\delta^{*}<\kappa_{0}$ that is a cutpoint in $\mathcal{M}^{(0)}|\kappa_{0}$, let
@@
be the result of a fully backgrounded $L[E]$-construction over $\emptyset$ in $\mathcal{M}^{(0)}|\delta_{1}$ using extenders with critical point above $\delta^{*}$, in other words, $\mathcal{L}_{\eta^{*}}=(L[E]^{\geq\delta^{*}}(\emptyset))^{\mathcal{M}^{(0)}|\delta_{1}}$. Let $\gamma\leq\eta^{*}$ be the least ordinal such that
@@
As $\kappa_{0}$ is a limit of Woodin cardinals in $\mathcal{M}^{(0)}$ (cf. [14, Lemma 2.14]) [^18] We will prove a weaker statement that still suffices to show that these constructions reach Woodin cardinals for the general case $i>0$ in the next subsection. and the least ${<}\delta_{1}$-strong cardinal in $\mathcal{W}$, there are such Woodin cardinals in $(Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma}))^{\mathcal{W}}$ cofinally below $\kappa_{0}$. In particular, for any cutpoint $\delta^{*}<\kappa_{0}$ in $\mathcal{M}^{(0)}$ and $\gamma\leq\eta^{*}$ as above, $\mathcal{L}_{\gamma}\cap\operatorname{Ord}<\kappa_{0}$. We start with arguing that these models $\mathcal{N}_{\delta^{*}}$ and $\mathcal{N}_{\delta^{*}}^{+}$ are in fact inside $\mathcal{M}^{(0)}$. This is a consequence of the following more general lemma, see also [14, Lemma 2.12].
@@
The function $X\mapsto(Lp^{\infty}(X))^{\mathcal{W}}$ for $X\in\mathcal{M}^{(0)}|\kappa_{0}$ is definable over $\mathcal{M}^{(0)}|\delta_{1}$.
@@
Let $\delta^{*}<\kappa_{0}$ be a cutpoint in $\mathcal{M}^{(0)}|\delta_{1}$ with $X\cap\operatorname{Ord}<\delta^{*}$ and let $\mathcal{L}_{\eta^{*}}=(L[E]^{\geq\delta^{*}}(X))^{\mathcal{M}^{(0)}|\delta_{1}}$ be the result of a fully backgrounded $L[E]$-construction over $X$ in $\mathcal{M}^{(0)}|\delta_{1}$ using extenders with critical point above $\delta^{*}$. Let
@@
Let $\eta>X\cap\operatorname{Ord}$ be a cardinal in $\mathcal{M}^{(0)}|\kappa_{0}$. Then
@@
The direction $\mathcal{O}^{\mathcal{L}_{\eta^{*}}}_{\eta}\unlhd(Lp^{\infty}(\mathcal{L}_{\eta}))^{\mathcal{W}}$ is clear. For the other direction, fix an arbitrary sound $\mathcal{N}\unlhd Lp^{\infty}(\mathcal{L}_{\eta})$ in $\mathcal{W}$ with $\mathcal{L}_{\eta}\unlhd\mathcal{N}$ and $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$. Let $E$ be the index-least extender on the extender sequence of $\mathcal{L}_{\eta^{*}}$ such that $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ is a cutpoint in $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$. It suffices to show that $\mathcal{N}\unlhd\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$. Note that both $\mathcal{N}$ and $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ are fully iterable in $\mathcal{W}$ above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ as the extenders used to construct $\mathcal{L}_{\eta^{*}}$ are fully backgrounded by extenders in $\mathcal{M}^{(0)}|\delta_{1}$ which are themselves fully backgrounded by extenders in $\mathcal{W}$. This uses the observation that the extenders in $\mathcal{M}^{(0)}$ with index below $\delta_{1}$ are all added by the fully backgrounded part of the $\mathcal{M}$-construction.
By universality of fully backgrounded constructions (cf. [13, Lemma 2.13], this uses that there are no inner models with a superstrong cardinal in $\mathcal{W}$), $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ wins the coiteration with $\mathcal{N}$. As $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$ and the coiteration is above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$, the $\mathcal{N}$-side cannot move as otherwise it would have to drop. Suppose the $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$-side moves to an iterate $\mathcal{W}^{*}$ of $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ such that $\mathcal{N}\unlhd\mathcal{W}^{*}$. If the iteration from $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ to $\mathcal{W}^{*}$ drops on the main branch, $\mathcal{W}^{*}$ is not fully sound, so $\mathcal{N}\lhd\mathcal{W}^{*}$. As $\rho_{\omega}(\mathcal{N})=\mathcal{L}_{\eta}\cap\operatorname{Ord}$, $\mathcal{N}$ is in fact a subset of $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ in $\mathcal{W}^{*}$. But iterations above $\mathcal{L}_{\eta}\cap\operatorname{Ord}$ do not add any new subsets of $\mathcal{L}_{\eta}\cap\operatorname{Ord}$, thus $\mathcal{N}\unlhd\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$, as desired. If the iteration from $\operatorname{Ult}(\mathcal{L}_{\eta^{*}},E)$ to $\mathcal{W}^{*}$ does not drop on the main branch, $\mathcal{W}^{*}\cap\operatorname{Ord}=\mathcal{L}_{\eta^{*}}\cap\operatorname{Ord}=\delta_{1}$ and again $\mathcal{N}\lhd\mathcal{W}^{*}$. So we can finish the argument as in the dropping case. ∎
The lemma now follows from the next claim, which is proved like Claim 2.13 in the proof of [14, Lemma 2.12]. It uses Clause (6) in Definition 3.20 and generic interpretability of the involved iteration strategies with branch condensation.
@@
For any $X\in\mathcal{M}^{(0)}|\kappa_{0}$, $\mathcal{N}\unlhd(Lp^{\infty}(X))^{\mathcal{W}}$ if, and only if, there is some $\mathcal{R}\unlhd\mathcal{O}^{\mathcal{L}_{\eta^{*}}}_{\eta}$ such that
@@
We start internalizing $\mathcal{P}_{\infty}^{0}$ into $\mathcal{M}^{(0)}$ with the following observation, that still takes place in $\mathcal{W}$.
@@
Let $\delta^{*}<\kappa_{0}$ be a cutpoint in $\mathcal{M}^{(0)}$. Then, in $\mathcal{W}$, $\mathcal{N}_{\delta^{*}}^{+}$ is a $\Sigma_{0}$-iterate of $\mathcal{P}^{0}$ and the iteration does not drop on the main branch.
@@
Work in $\mathcal{W}$ and compare $\mathcal{M}^{(0)}|\delta_{0}$ against the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$ giving rise to $\mathcal{N}_{\delta^{*}}$. By stationarity and universality of fully backgrounded constructions [^19] Cf. [13, Lemmas 2.11 and 2.13], note that these results also hold in our setting where $\mathcal{N}_{\delta^{*}}$ is the result of a fully backgrounded construction in $\mathcal{M}^{(0)}|\delta_{1}$, that itself is an initial segment of the result of a fully backgrounded construction in $\mathcal{W}$., $\mathcal{M}^{(0)}|\delta_{0}$ iterates via $\Sigma_{0}$ without drops on the main branch to some level of the construction $(\mathcal{L}_{\eta}\mid\eta\leq\eta^{*})$. As $\mathcal{N}_{\delta^{*}}^{+}=Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma})$ for $\gamma$ the least ordinal such that $\mathcal{L}_{\gamma}\cap\operatorname{Ord}$ is Woodin in $Lp_{\omega}^{\infty}(\mathcal{L}_{\gamma})$, it follows by super fullness preservation of $\Sigma_{0}$ that in fact $\mathcal{P}^{0}=Lp_{\omega}^{\infty}(\mathcal{W}|\delta_{0})$ iterates via $\Sigma_{0}$ to $\mathcal{N}_{\delta^{*}}^{+}$. ∎